# Rethinking the Role of Prompting Strategies in LLM Test-Time Scaling: A Perspective of Probability Theory

### v1

@@ Appendix B Proofs
Appendix C Detailed Results
Here we display the scaling performance of different prompting strategies on each LLM and benchmark under given sampling time $N$ and cost $O$, as shown in Figures 8 to 15. We find that, aside from CoT, DiP also exhibits superior performance compared to other complex prompting strategies on certain models and datasets, e.g., Gemini-1.5-Flash on MATH. This also comes from the two reasons, i.e., DiP has more hard questions and easy questions and a flat probability distribution of wrong answers on the specific dataset. This phenomenon is particularly prominent on powerful LLMs such as Gemini-1.5-Flash on GSM8K and GSM-Hard, where DiP and CoT exhibit comparable performance. Almost 83% of results satisfy that CoT or DiP performs best as significantly scaling. Besides, this trend is also observed on other prompting strategies on few datasets and models. This encourages us to fully unleash the potential of simple prompting strategies, and indicates that the scaling performance does not only depend on the prompting strategies’ pass@1 accuracy.
@@ Appendix D More Discussions on Improving the Scaling Performance
In this section, we will discuss more about our further exploration on the two ways to improve the scaling performance. We display more results of (1) adaptively scaling based on the question difficulty, (2) dynamically choosing the optimal $\mathbf{P}_{i}$ and (3) combining adaptively scaling and dynamically choosing the optimal $\mathbf{P}_{i}$ in Section D.1, D.2 and D.3, respectively.
@@ D.1 Adaptively Scaling Based on the Difficulty
D.2 Dynamically Choosing the Optimal $\mathbf{P}_{i}$
Figures 21 to 25 display the results on GSM8K on LLaMA-3-8B-Instruct, GLM-4-9B-Chat, Phi-3.5-mini-Instruct, Gemini-1.5-Flash and GPT-4o-mini, respectively. It can be observed that all LLMs tend to believe that CoT is the best prompting strategy, while CoT does not excel at every question. With oracles to provide the optimal $\mathbf{P}_{i}$ labels, all LLMs demonstrate significant performance improvements, even only with one sampling time, proving the enormous potential of this method. we will explore how to approach this upper bound in the future.
@@ D.3 Combining Adaptively Scaling and Dynamically Choosing the Optimal $\mathbf{P}_{i}$
Appendix E Implementation Details and Prompts
We use vllm (Kwon et al., 2023) to deploy open-sourced LLMs, with top-p = 0.9 and temperature = 0.7. For close-sourced LLMs, we use their APIs with default settings. We set the content safety detection threshold of Gemini-1.5-Flash to zero to prevent erroneous judgments that may result in null outputs.
Following (Wang et al., 2024; Lightman et al., 2024; Qi et al., 2025), we use MATH-500, a subset of representative problems from the MATH dataset to speed up the evaluation. We use the test split of each dataset. The license for all datasets is CC-BY 4.0 or others for open academic research. The number of samples on each dataset is shown in Table 5. We ensure our use of existing artifacts is aligned with their intended purposes. All of them are public English datasets for academic research. On GSM8K and GSM-Hard, we use the same 1-shot prompt in the original paper of Least-to-Most (Zhou et al., 2023) shown in Figure 39 and Figure 40. On other datasets, we use the 0-shot prompt shown in Figure 41. We use the same prompt in Analogous Prompting (Yasunaga et al., 2024), and guide the LLM to recall one analogous problem. We use the same 1-shot prompt in Step-Back Prompting (Zheng et al., 2024) on MMLU, and apply their prompt designed for reasoning tasks on other datasets. We use the same prompt in 0-shot Chain-of-Thought Kojima et al. (2022), Multi-Agent Debate Du et al. (2024) and Self-Refine Huang et al. (2024) on all datasets. The prompts are shown in Figures 37 to 46.

### v2

@@ Appendix D More Discussions on Improving the Scaling Performance
In this section, we will discuss more about our further exploration of the two ways to improve the scaling performance. We display more results of (1) adaptively scaling based on the question difficulty, (2) dynamically choosing the optimal $\mathbf{P}_{i}$ and (3) combining adaptively scaling and dynamically choosing the optimal $\mathbf{P}_{i}$ in Section D.1, D.2 and D.3, respectively.
@@ D.1 Adaptively Scaling Based on the Difficulty
D.2 Dynamically Choosing the Optimal $\mathbf{P}_{i}$
Figures 21 to 25 display the results on GSM8K on LLaMA-3-8B-Instruct, GLM-4-9B-Chat, Phi-3.5-mini-Instruct, Gemini-1.5-Flash and GPT-4o-mini, respectively. It can be observed that all LLMs tend to believe that CoT is the best prompting strategy, while CoT does not excel at every question. With oracles to provide the optimal $\mathbf{P}_{i}$ labels, all LLMs demonstrate significant performance improvements, even with only one sampling time, proving the enormous potential of this method. we will explore how to approach this upper bound in the future.
@@ D.3 Combining Adaptively Scaling and Dynamically Choosing the Optimal $\mathbf{P}_{i}$
Appendix E Extended Experiments
We extend experiments with Qwen2.5-7B-Instruct on two more challenging reasoning benchmarks, GPQA (Rein et al., 2024) and AIME2024 (Mathematical Association of America, 2024), further verifying the generality of our findings and methods. Figure 37 displays the accuracy of each $\mathbf{P}_{i}$ with Qwen2.5-7B-Instruct on GQPA and AIME. Simple CoT and DiP gradually dominate as scaling sampling times on GPQA and AIME, respectively. On GPQA, we can see that the performance fluctuates as scaling, rather than constantly increasing. This can be attributed to the ratio of easy/hard questions and their accuracies in the entire dataset. Given AnP as an example, Figure 39 reports its scaling performance on the easy/hard question subsets of GPQA. Performance on easy questions monotonically improves with increasing sampling times, while accuracy on hard questions exhibits a corresponding decline. This fundamental trade-off induces characteristic oscillation in the aggregate accuracy, with consistent replication across all tested prompting strategies, substantiating our theoretical framework.
Figure 38 shows the results of adaptively scaling on each $\mathbf{P}_{i}$. “Adaptive” approach demonstrates evident performance gains over “Vanilla” baselines across CoT, L2M, SBP, and AnP, indicating the model’s intrinsic capability to assess question difficulty. Under “Oracle” conditions, it can achieve further performance amplification.
Figure 40 reports the results of dynamically choosing the optimal $\mathbf{P}_{i}$ on Qwen2.5-7B-Instruct on GPQA. “Dynamic” approach achieves median-range performance across all tested $\mathbf{P}_{i}$, quantitatively confirming the model’s suboptimal strategy selection capacity. Strikingly, “Oracle” intervention enables dramatic performance elevation, with 65.4% accuracy at $N=1$.
Figure 41 summarizes the results of combining adaptively scaling and dynamically choosing the optimal $\mathbf{P}_{i}$ on Qwen2.5-7B-Instrcut on GPQA, which further enormously enhances the scaling performance, with 75.7% accuracy at $N=100$.
Figure 37: Accuracy of each $\mathbf{P}_{i}$ with Qwen2.5-7B-Instruct on GQPA and AIME2024.
Figure 38: Results of adaptively scaling based on the question difficulty on Qwen2.5-7B-Instruct on GPQA.
Figure 39: Scaling performance of AnP on the easy/hard question subsets of GPQA. The accuracy on easy questions is non-decreasing with the sampling time, while it exhibits a general declining trend on hard questions. This also holds for all other prompting strategies.
Figure 40: Results of dynamically choosing the optimal $\mathbf{P}_{i}$ on Qwen2.5-7B-Instruct on GPQA.
Figure 41: Results of combining adaptively scaling and dynamically choosing the optimal $\mathbf{P}_{i}$ on Qwen2.5-7B-Instrcut on GPQA.
Appendix F Implementation Details and Prompts
We use vllm (Kwon et al., 2023) to deploy open-sourced LLMs, with top-p = 0.9 and temperature = 0.7. For closed-sourced LLMs, we use their APIs with default settings. We set the content safety detection threshold of Gemini-1.5-Flash to zero to prevent erroneous judgments that may result in null outputs.
Following (Wang et al., 2024; Lightman et al., 2024; Qi et al., 2025), we use MATH-500, a subset of representative problems from the MATH dataset to speed up the evaluation. We use the test split of each dataset. The license for all datasets is CC-BY 4.0 or others for open academic research. The number of samples on each dataset is shown in Table 5. We ensure our use of existing artifacts is aligned with their intended purposes. All of them are public English datasets for academic research. On GSM8K and GSM-Hard, we use the same 1-shot prompt in the original paper of Least-to-Most (Zhou et al., 2023) shown in Figure 44 and Figure 45. On other datasets, we use the 0-shot prompt shown in Figure 46. We use the same prompt in Analogous Prompting (Yasunaga et al., 2024), and guide the LLM to recall one analogous problem. We use the same 1-shot prompt in Step-Back Prompting (Zheng et al., 2024) on MMLU, and apply their prompt designed for reasoning tasks on other datasets. We use the same prompt in 0-shot Chain-of-Thought Kojima et al. (2022), Multi-Agent Debate Du et al. (2024) and Self-Refine Huang et al. (2024) on all datasets. The prompts are shown in Figures 42 to 51.